OpenStudy (anonymous):

Copying machines A and B, working together, can prepare enough copies of the annual report for the board of directors meeting in 2 hours. Machine A, working alone, requires 3 hours to do the job. How long would it take Machine B to do the job by itself?

OpenStudy (anonymous):

Let B be the number of hours for B to do the job alone. Solve the following for B: \[\frac{1}{B}+\frac{1}{3}=\frac{1}{2} \]B = 6 hours